LATEX(1) LATEX(1) NAME latex−structuredtextformattingandtypesetting SYNOPSIS latex[firstline] DESCRIPTION TheLaTeXlanguageisdescribedinLaTeX−ADocumentPreparationSystem,byLeslieLamport, publishedbyAddison-Wesley.LaTeXisaTeXmacropackage,notamodificationtotheTeX sourceprogram,soallthecapabilitiesdescribedintex(1)arepresent. TheLaTeXmacrosencouragewriterstothinkaboutthecontentoftheirdocuments,ratherthan theform.Theideal,notalwaysrealized,istohavenoformattingcommands(like‘‘switchto italic’’or‘‘skip2picas’’)inthedocumentatall;instead,everythingisdonebyspecificmarkup instructions:‘‘emphasize’’,‘‘startasection’’. SEEALSO slitex(1),tex(1) 7Jan92 1